# Quantum-Enhanced Generative Adversarial Networks: Comparative Analysis of Classical and Hybrid Quantum–Classical Generative Adversarial Networks

### v1

@@ III-A5 Training Strategy and Metrics
III-B Authors and Affiliations
The class file is designed for, but not limited to, six authors. A minimum of one author is required for all conference articles. Author names should be listed starting from left to right and then moving down to the next line. This is the author sequence that will be used in future citations and by indexing services. Names should not be listed in columns nor group by affiliation. Please keep your affiliations as succinct as possible (for example, do not differentiate among departments of the same organization).
III-C Quantum Circuit Visualisation
@@
III-C1 Quantum Circuit Diagram
@@
III-C2 Noisy Density Matrix (Cityscape)
@@
III-C3 Pauli Vector Representation
@@ III-C4 Bloch Multivector
III-C5 Noisy Bitstring Count Histogram
@@
III-D Generator and Discriminator Training Dynamics
III-D1 Discriminator Loss Dynamics
@@
III-D2 Generator Loss Dynamics
@@
As illustrated in Fig. 7, the Classical GAN $\mathcal{L}_{G}$ steadily decreases when $N\approx 0.7M$, reaching values of $\mathcal{L}_{G}\approx 3$. Although this suggests that the generator is improving, the concurrent increase in $\mathcal{L}_{D}$ (from Section III-D1) suggests that this improvement may be artificial. The possibility of an overpowered generator dominating a weakening discriminator resulting in Mode Collapse cannot be ruled out in this scenario.
@@
III-E Combined Interpretation of Training Dynamics
@@ III-F FID and KID Evaluation
III-F1 Fréchet Inception Distance (FID)
@@
III-F2 Kernel Inception Distance (KID)
@@
III-F3 Evaluation Metrics Discussion

### v2

@@ III-A5 Training Strategy and Metrics
III-B Quantum Circuit Visualisation
@@
III-B1 Quantum Circuit Diagram
@@
III-B2 Noisy Density Matrix (Cityscape)
@@
III-B3 Pauli Vector Representation
@@ III-B4 Bloch Multivector
III-B5 Noisy Bitstring Count Histogram
@@
III-C Generator and Discriminator Training Dynamics
III-C1 Discriminator Loss Dynamics
@@
III-C2 Generator Loss Dynamics
@@
As illustrated in Fig. 7, the Classical GAN $\mathcal{L}_{G}$ steadily decreases when $N\approx 0.7M$, reaching values of $\mathcal{L}_{G}\approx 3$. Although this suggests that the generator is improving, the concurrent increase in $\mathcal{L}_{D}$ (from Section III-C1) suggests that this improvement may be artificial. The possibility of an overpowered generator dominating a weakening discriminator resulting in Mode Collapse cannot be ruled out in this scenario.
@@
III-D Combined Interpretation of Training Dynamics
@@ III-E FID and KID Evaluation
III-E1 Fréchet Inception Distance (FID)
@@
III-E2 Kernel Inception Distance (KID)
@@
III-E3 Evaluation Metrics Discussion